3+(x)is differentiable or not here ()is mod

Grade:12

1 Answers

naga Vishnu
10 years ago

3+(x) is differntiable for all real values of x except for x=0

      The reson for this is

                              When x<0 then 3+(x) = 3-x

                                              then Derivative of 3+(x) is -1

                              When x>0 then 3+(x) = 3+x

                                              then Derivative of 3+(x) is 1

             From this we can see that at x=0 the Right side derivatie of 3+(x) is not equal to the left side derivative.

                               Hence 3+(x) is not differentiable at x=0.